OpenStudy (anonymous):

If the measures of three angles in a triangle are (2x+15)º, (10x – 20)º, and (3x + 2)º, then find the value of 15x – 3. I am getting 180. Is it right ?

OpenStudy (saifoo.khan):

The sum of all the internal angles of a triangle = 180*

OpenStudy (anonymous):

(2x+10)+10x-20)+(3x+2)=180 15x-3=180 15x=183 x=183/15=12.2 x=12.2 (15x-3)=15(12.2)-3=180

OpenStudy (anonymous):

Is it right

OpenStudy (saifoo.khan):

I guess you did a mistake? 2x+15

OpenStudy (anonymous):

Please can you check it

OpenStudy (saifoo.khan):

it's 2x+15!

OpenStudy (anonymous):

I got 15x -3 = 180

OpenStudy (anonymous):

I wrote wrong here. but I put the correct numbers

OpenStudy (anonymous):

yes I got the same

OpenStudy (anonymous):

2x + 15 + 10x – 20 + 3x + 2 = 180

OpenStudy (saifoo.khan):

x = 12.2 is right!

OpenStudy (anonymous):

My textbook saying 90 as a answer